# PHYSICISTS JOKE

### v1

@@ Abstract
This work presents a reconstructed English edition of the seminal humor anthology Physicists Joke (Fiziki Shutyat), originally published in the USSR in 1966. While the Soviet edition served as a significant cultural artifact of the 1960s, a large portion of its content consisted of Russian translations of articles originally written in English. To avoid the loss of nuance inherent in back-translation, this edition does not simply translate the Russian text. Instead, the original English source materials have been identified and restored from archives of journals such as Physics Today, The Journal of Irreproducible Results, New Scientist, and Proceedings of the IEEE. The specific structural framework of the 1966 book has been preserved, along with English translations of the unique Russian scientific folklore, anecdotes, and editorial interludes compiled by the original Soviet editors. This collection offers a glimpse into the sociology of science and the shared professional culture of the "Golden Age" of mid-20th-century physics.
@@
Note on the 2025 English Translation
This book, originally published in the USSR in 1966, was largely a collection of humorous scientific articles translated from English into Russian. To ensure the highest fidelity in this English edition, I have tried not to simply translate the Russian text back into English, which would result in a "broken telephone" loss of nuance.
Instead, wherever possible, I have located the original English source texts (from journals such as Physics Today, The Journal of Irreproducible Results, New Scientist, etc.) and restored the authentic wording used by the original authors. For the short anecdotes, folklore, and specific parodies where the original English text was unavailable or where the Russian version differed significantly from the source, a direct translation from the Russian text has been performed, striving to preserve the humor and style of the 1966 edition.
I would like to acknowledge the assistance of the AI model Google Gemini 3 Pro in locating the original English source texts and translating the specific Russian scientific folklore contained in this volume.
x-Professor Igor Halperin, December 2025
@@ 1.2 Confessions of an Acoustical Engineer
When I was an assistant, I worked like a horse, and earned so little that I barely had enough for food. My boss always "quipped" about my mental abilities and spent half the day explaining to me what was perfectly clear without him, and then was surprised and threw up his hands that the work was not yet done. He assigned me to develop drawings of unfeasible designs invented by thinkers in our laboratory who were hovering in the clouds. I had to finish and refine everything for everyone to force these constructions to work at least somehow. When I came to the boss with some clever solution, he put it aside and said that this was not what they wanted… Sometimes I worked on a similar project for a year and came to him literally with a masterpiece. Then he would declare: "Very good, my boy, but management decided to tackle a somewhat different topic." Consequently, back to the drawing board…
By that time, as I became a group leader, the situation in the institute had changed. Assistants completely let themselves go and did not know how to do anything, but were offended by every remark. Suppose a trifling job needed to be done. I would have done it myself in a couple of hours. But I had to spend half the day explaining to my assistants why this work should be done at all, why it should be done faster than any other, why it should be done this way and not otherwise, and why necessarily by a certain deadline. When these deadlines passed, I had to listen to half a day of incoherent explanations why the work was not yet done* why nothing works (and, perhaps, will not work) and why we should start everything from the beginning and do everything "as needed". But this "needed" path turned out to be so confused that it usually took at least a year to figure it out, and even less than a year to get some results. At this stage of work somehow, with the help of diplomacy and flattery, persuasion and requests, it was possible to force assistants to make ends meet and push this work off their necks.
Even worse was the case with my administrative duties. I was forced to write so many reports and proposals that this alone would take all my working time. But I still had to answer all current correspondence, answer all phone calls, and receive visitors who came to me in a continuous stream. I had to hire new people and at the same time take care that the old ones remained satisfied with it. It was also assumed that I must take care of the representativeness of the company and for this purpose attend professional meetings, give work in committees, and arrange seminars.
In the end, I started a small, but my own business. Unfortunately, I was unlucky with technical managers and administrators. They have changed much for the worse. They do not strive at all to ensure that something is done by their subordinates. They organize the matter in such a way that everything in the world collapses, and then reorganization begins. They travel, converse, arrange seminars, attend meetings, participate in technical commissions, that is, do anything, but only do not work for the company. The intellectual power enclosed in them and their assistants is colossal, but it is spent not on creation, but on destruction. If the Nobel Prize were awarded for excuses, then our laboratory would have received it long ago.
Now, having gained bitter experience, I dream of becoming an assistant again. An assistant still lives easier. But, unfortunately, I am already married and cannot afford this luxury.
@@ 1.3 A Contribution to the Mathematical Theory of Big Game Hunting
Author: H. Pétard
Original Source: The American Mathematical Monthly, Vol. 45, No. 7 (1938).
@@
    The Inversive Geometry Method. We place a spherical cage in the desert, enter it, and lock it. We perform an inversion with respect to the cage. The lion is then in the cage, and we are outside.
@@
    The Method of Projective Geometry. Without loss of generality, we can regard the Sahara Desert as a plane. Project the plane into a line, and then project the line into a point inside the cage. The lion is projected into the same point.
@@
    The Bolzano-Weierstrass Method. Bisect the desert by a line running N-S. The lion is either in the East or West section. Bisect that section E-W. The lion is in one of the two parts. Continue this process indefinitely…
@@ 1.5 Earth as a Man-Controlled Spaceship
And now let’s look into a more distant future. Personally, I generally cannot understand why astronauts dream of getting into interstellar space. In a rocket, it will be terribly cramped. And with food, they will have to cut themselves down significantly. But this is still half the trouble. The main nuisance is that the astronaut in the rocket will be in the same position as a man placed against a beam of fast protons coming out of a powerful accelerator (see figure). I feel very sorry for the poor astronaut; about his sad fate, I even composed a ballad*:
Ballad of an Astronaut
From the beta-inverter and gamma-converter
Remained only the plating.
And the ion cannon, like an empty party popper,
Sticks out, good for nothing.
All mesons decayed, all neutrons decayed,
All visible light radiated away.
According to Coulomb’s law, protons scattered,
For leptons, there is no hope.
The damaged reactor rattles like a tractor,
In the biochamber – rot and mold.
The nozzle is clogged, the bottom is leaking,
And vacuum lashes into the gap…
He flew to Orion, but a flux of gravitons
Crossed his path unexpectedly.
Deviating from course and exhausting all resources,
He barely managed to slip away.
Made a huge hook, flew around half the Universe
And now on an empty ship
In a straight line he was returning home,
Approaching the planet Earth.
But fighting with gravity, super-super-super-acceleration,
He slowed down the hands of the clock.
And the hands froze, on Earth passed
Thousands of thousands of centuries.
Here are the native planets… God! Is this the Sun?
Dark red, slightly warm ball…
Above the Earth it smokes,
Above the Earth swirls Hydrogen, cold steam.
What is this?
Where is the human tribe? – In unknown, distant worlds.
Their children grow
Already on a new planet, And Earth is all in cosmic ice.
Swearing and crying
From such a failure, the Astronaut turned the lever,
And B rang out
And A rang out
And X rang out –
BANG!
* Free translation from English by V. Turchin.
@@ 1.7 Typology in Scientific Research
    THE DISCOVERER (The Idea Man). These scientists are always "giving out" new ideas. Their brains are always ready to sink their teeth into random prey. Good scientific training allows them to quickly assess the importance of an observed fact and formulate an idea, after which the hypothesis is ready (or at least a working hypothesis). Then they either verify it experimentally themselves or leave it to others to worry about, deriving pleasure from the purely speculative solution of the problem.
@@
    THE EXPLOITER. These are researchers with a quick grip; their eyes and ears are constantly open. Such a scientist is rarely found in his own laboratory; he prefers to spend his time in discussions with colleagues from other laboratories and institutes, especially if these colleagues are working on something that interests him. He never lacks good ideas, which, although they were not born in his head, turn into interesting articles, generously sprinkled with references to "private communications."
@@
    THE CONNOISSEUR (The Critic). The mental abilities of such a man far exceed his capabilities (and desire) to conduct his own experiments. He is capable of evaluating (and does evaluate) good work, and often does this better than the author himself. A critical mind, combined with innate inconsistency, is the reason why the results of each subsequent series of measurements significantly differ from all previous ones; this prevents such a scientist from publishing anything unless he has a decisive boss.
@@
    THE IMPROVER. He resembles the "Connoisseur," but possesses slightly higher productivity. His achievements are represented by very few, but excellent articles, based on experiments which he repeated so many times that all unexpected or unpredicted results managed to be discarded with the help of sophisticated statistical processing.
@@
    THE MAN ON THE LEVEL (The Well-Read Man). He knows everything that is worth knowing. Unlike the "Exploiter," he spends all his time in the library, where it is rare for anyone to beat him to receiving the latest issue of a journal.
@@
    THE CO-AUTHOR. This type has perfectly mastered the art of scientific diplomacy. He achieves the inclusion of his name in the lists of authors of most articles published by employees of the department where he works, without any pressure, and his contribution is sometimes expressed merely in deciding a question—whether to use the conjunction "and" in the title of the article. Some people hold the opinion that the "Co-author" is almost the same as the "Advisor," but what an advisor is, everyone who plays cards or chess knows.
@@
    THE ADVISOR (The Consultant). Advisors, who are found most often in institutions engaged in fundamental research, should not be confused with consultants who, engaged in science, give advice to the relevant branch of industry, giving people appropriate advice for appropriate remuneration.
@@
    THE GADGETEER. In the world of modern science, there is research, the conduct of which is absolutely impossible without a solid set of instruments. Opportunities that were available to Archimedes and Newton are now limited to no one (except, perhaps, theoretical physicists), and a well-equipped laboratory is as necessary for productive research as a typewriter is for writing reports to management. Therefore, some scientific workers consider the meaning of their life to be obtaining and (if possible) using the largest possible number of extremely modern instruments. Visiting such a laboratory is a sheer pleasure. The soul simply rejoices at the sight of rooms packed with ultra-modern equipment that sparkles with glass and nickel; awe-inspiring names of numerous installations are used more to impress visitors than for any other purpose.
@@
    THE PUBLISHER. By this term, for lack of a better one, we will denote those who, by any numerous means at an exponentially increasing rate, lengthen the list of their scientific works. A variation of this type is:
    • The Re-hasher (The Ruminant) (a name associated with the peculiarities of the digestive process in certain mammals). Such a person usually settles in some underdeveloped country. He publishes his observations there and findings, prefacing them with the introduction: "For the first time in history… (insert name of country) it was observed…", after which a translation into the local language of some work done by other people in another place is honestly reproduced.
    Another representative of the same type:
    • The Multiplier. An individual who lays out the results of his work or his speculative reasoning into the largest possible number of small packets with labels "Article," "Letter to the Editor," "Short Communication," etc., and successfully increases the list of his works in this way.
@@
    THE CORRESPONDENT. To get acquainted with this type of scientist, we refer the reader to the last pages of any scientific journal, where headings such as "Notes," "Brief Communications," "Preliminary Results," or "Letters to the Editor" are dazzling. There, every Correspondent reports something truly important, very similar to a great discovery, which follows to publish as quickly as possible before someone else does… Such messages end with the words: "A detailed description of experiments (or results) will be published in such-and-such an edition (or in the near future)." In 50% of cases, the promised publication never appears, since the results of repeated experiments beat off any interest in the idea itself.

### v2

# PHYSICISTS ARE STILL JOKING Almanac of Physics and Science Humor (1966–2025) Containing the restored translations of the Russian classics:

@@ Abstract
This volume, Physicists Are Still Joking, serves as a definitive almanac of scientific humor spanning sixty years. It traces the evolution of professional folklore across geopolitical divides and technological eras. Part I restores the classic 1966 anthology Physicists Joke, which originally served as a window for Soviet scientists into the best traditions of Western scientific humor; it consists primarily of articles translated from English, here meticulously restored to their original wording. Part II presents the 1992 sequel, Physicists Keep Joking, which captures the shift toward an original, introspective Russian scientific folklore born during the end of the Cold War and the collapse of the Soviet Union. Part III: Still Joking explores the modern digital age, compiling contemporary science humor from physics, astronomy, biology, computer science and AI research. While the tools of science have evolved from slide rules to neural networks, the tradition of skeptical, self-referential wit remains a constant. Spanning from the "Golden Age" of vacuum tubes to the era of AI and Large Language Models, this collection documents the enduring ability of scientists to laugh at the universe and themselves.
Preface to the First Edition
The genesis of this almanac lies in an attempt to preserve and restore a unique cultural artifact of the Cold War era: the Soviet anthology Physicists Joke (Физики шутят), originally published in 1966.
The 1966 collection was a cultural phenomenon in the USSR. It offered a glimpse into the sociology of the "Golden Age" of physics—a time of cyclotrons, slide rules, and unbounded optimism. However, it is important to note that the majority of that book consisted of Russian translations of articles originally written in English and German. It was, in effect, a curation of the best Western scientific humor, seen through the eyes of Soviet physicists behind the Iron Curtain.
In preparing the initial draft of this edition, I realized that simply translating the Russian text back into English would result in a "broken telephone" effect, where the nuance and wit of the originals would be lost in double translation. Therefore, the primary focus of the first edition was linguistic archaeology. Wherever possible, I located the original English source texts—scouring archives of Physics Today, The Journal of Irreproducible Results, New Scientist, and conference proceedings from the 1950s and 60s—to restore the authentic wording used by the original authors.
For the unique Russian scientific folklore, anecdotes, and editorial interludes where no English original existed, a direct translation has been performed, striving to preserve the specific voice of the Soviet physicist.
I would like to acknowledge the assistance of the AI model Google Gemini 3 Pro in locating these archival texts and assisting with the translation of Russian folklore.
Igor Halperin
December 15, 2025
Preface to the Second Edition
Physics does not stand still, and neither does the humor of physicists.
This Second Edition transforms the work from a restoration project into a comprehensive almanac. It reflects the realization that scientific humor is not static; it evolves to adapt to new geopolitical realities and new technologies, while maintaining the traditions of the past.
Part II: Resilience in the Face of a Collapse (1992)
While the 1966 book was largely a mirror reflecting Western humor into the Soviet Union, the 1992 sequel, Physicists Keep Joking (Физики всё ещё шутят), marked a distinct shift. Published right after the collapse of the USSR, it ceased looking outward for material and instead captured the unique, original voice of the Russian scientific community.
The humor in Part II is sharper, more cynical, and deeply intertwined with the specific hardships of the time—the bureaucracy, the funding crises, and the crumbling of the "ivory tower." It documents how the scientific tradition survived the end of the Cold War through irony and resilience.
Part III: Science Humor in the Digital Age (1990s–Present)
Part III brings the almanac into the present day. Both modern physics and modern science in general are vastly different from where they stood in the 1990s. The main focal points of physicists’ attention have shifted—from the excitement over string theory and the search for the Higgs boson to gravitational wave astronomy, quantum computing, and the mysteries of dark energy. The tools of the trade have evolved just as dramatically, from early workstations and FORTRAN code to GPU clusters, machine learning pipelines, and Large Language Models.
Yet perhaps an even more profound transformation since the 1990s has been the rise of interdisciplinarity. Modern science increasingly dissolves the boundaries between traditional fields. Physicists today actively collaborate with pure and applied mathematicians, chemists, biologists, computer scientists, and AI researchers. A condensed matter physicist might work alongside a machine learning specialist; a cosmologist might co-author papers with statisticians and data scientists; a biophysicist might find themselves debugging Python code with a computational neuroscientist.
Inevitably, these increasing interdisciplinary interactions have led to a cross-pollination of scientific humor traditions as well. The jokes, parodies, and satirical papers now travel freely across departmental boundaries. Physicists have adopted the self-deprecating wit of mathematicians; computer scientists have inherited the experimentalist-vs-theorist banter; and the AI community has developed its own rich folklore that borrows liberally from all its parent disciplines. This new section reflects that broader landscape.
To compile Part III, I undertook an extensive search of the modern digital landscape—scouring web archives that collect April Fools arXiv papers, the SIGBOVIK proceedings, and other online repositories of scientific humor. I looked beyond pure physics to include the folklore of mathematicians, biologists, computer scientists, and the AI community. The result is a collection that covers the era of string theory, big data, data science, and artificial intelligence. It shows that while the tools of science have changed from vacuum tubes and slide rules to neural networks and LLMs, the spirit of the scientist—skeptical, irreverent, and always ready to laugh at the universe and at themselves—remains wonderfully unchanged.
Acknowledgments
Since the release of the preliminary draft of this book, I have received a number of warm comments from scientists who grew up in the later USSR or former Soviet republics, where these books still hold a legendary status. Many people have shared memories of spending hours during their high school and university years reading and laughing at these collections.
It is a testament to the quality of the original selection that quotes from the 1966 edition are still popular to this day. One particular favorite, which I was reminded was especially popular in theory departments as a running joke about experimentalists (a convenient explanation for any "groundbreaking" anomaly in the data), remains:
"…And that was Wilkins hitting the seismograph!"
(…А это Уилкинс ударил по сейсмографу!)
I am deeply grateful to the readers of the first edition for their interest, appreciation, and for providing links to new materials that helped shape Part III. I would like to thank Alexander Barzykin, Arthur Bird, Elena Nazvanova, Dmitri Offengenden, and Mihail Turlakov for their valuable comments, suggestions and encouragement.
Igor Halperin
December 24, 2025
Part I Physicists Joke (1966)
@@ 1.2 Confessions of an Acoustical Engineer
When I was an assistant, I worked like a dog and earned barely enough to eat. My boss always made cracks about my mental abilities and spent half the day explaining things that were perfectly clear, and then threw up his hands in surprise that the work wasn’t done. He assigned me the task of drawing up blueprints for unworkable designs invented by the "thinkers" in our lab who had their heads in the clouds. I had to finish and refine everything to force these contraptions to work at least somehow.
By the time I became a Group Leader, the situation had changed. The assistants had let themselves go completely; they didn’t know how to do anything but took offense at every remark. Suppose a trifling job needed to be done. I could have done it myself in two hours. But I had to spend half a day explaining to my assistants why this work should be done at all, why it should be done this way and not otherwise.
Even worse were my administrative duties. I was forced to write so many reports and proposals that this alone took all my time. But I still had to answer correspondence, take phone calls, and receive visitors. It was assumed that I must take care of the "image" of the company and attend professional meetings, serve on committees, and arrange seminars.
In the end, I started a small business of my own. Unfortunately, I have had bad luck with engineering managers and administrators. They have changed much for the worse. They do not strive to get things done. They organize matters so that everything collapses, and then "reorganization" begins. They travel, converse, arrange seminars—they do anything but work. If the Nobel Prize were awarded for excuses, our laboratory would have received it long ago.
Now, having gained bitter experience, I dream of becoming an assistant again. An assistant still lives an easier life. But, unfortunately, I am already married and cannot afford the luxury.
@@ 1.3 A Contribution to the Mathematical Theory of Big Game Hunting
Author: H. Pétard, Princeton, New Jersey
Original Source: The American Mathematical Monthly, Vol. 45, No. 7 (Aug.–Sep., 1938).
1. Mathematical Methods
@@
    The Hilbert, or Axiomatic, Method. We place a locked cage at a given point of the desert. We then introduce the following logical system: Axiom I: The class of lions in the Sahara Desert is non-void. Axiom II: If there is a lion in the Sahara Desert, there is a lion in the cage. Rule of Procedure: If $p$ is a theorem, and "$p$ implies qq" is a theorem, then $q$ is a theorem. Theorem I: There is a lion in the cage.
@@
    The Method of Inversive Geometry. We place a spherical cage in the desert, enter it, and lock it. We perform an inversion with respect to the cage. The lion is then in the interior of the cage, and we are outside.
@@
    The Method of Projective Geometry. Without loss of generality, we may regard the Sahara Desert as a plane. Project the plane into a line, and then project the line into a point inside the cage. The lion is projected into the same point.
  4. 4.
    The Bolzano-Weierstrass Method. Bisect the desert by a line running N-S. The lion is either in the East or West section. Bisect that section by a line running E-W. The lion is in one of the two parts. We continue this process indefinitely, constructing a sufficiently strong fence at each step. The diameter of the chosen partitions approaches zero, so that the lion is ultimately surrounded by a fence of arbitrarily small perimeter.
  5. 5.
    The "Mengentheoretisch" Method. We observe that the desert is a separable space. It therefore contains an enumerable dense set of points, from which can be extracted a sequence having the lion as a limit. We then approach the lion stealthily along this sequence, bearing with us suitable equipment.
  6. 6.
    The Peano Method. Construct a continuous curve passing through every point of the desert. It has been remarked that it is possible to traverse such a curve in an arbitrarily short time. Armed with a spear, we traverse the curve in a time shorter than that in which a lion can move his own length.
  7. 7.
    A Topological Method. We observe that a lion has at least the connectivity of the torus. We transport the desert into four-space. It is then possible to carry out such a deformation that the lion can be returned to three-space in a knotted condition. He is then helpless.
  8. 8.
    The Cauchy, or Function Theoretic, Method. We consider an analytic lion-valued function $f(z)$. Let $\zeta$ be the cage. Consider the integral:
    $\frac{1}{2\pi i}\int_{C}\frac{f(z)}{z-\zeta}dz$
    where $C$ is the boundary of the desert; its value is $f(\zeta)$, i.e., a lion in the cage.
2. Methods from Theoretical Physics
  1. 1.
    The Dirac Method. We observe that wild lions are, ipso facto, not observable in the Sahara Desert. Consequently, if there are any lions in the Sahara, they are tame. The capture of a tame lion is left as an exercise for the reader.
  2. 2.
    The Schrödinger Method. At any given moment there is a non-zero probability that the lion is in the cage. Sit down and wait.
  3. 3.
    The Method of Nuclear Physics. Place a tame lion into the cage, and apply a Majorana exchange operator between it and a wild lion. Variant: Let us assume we would like to catch a lion ($L$) but we catch a lioness ($L^{\prime}$). We then place the latter in the cage and apply the Heisenberg exchange operator which exchanges the spins.
3. Methods from Experimental Physics
  1. 1.
    The Thermodynamical Method. We construct a semi-permeable membrane which allows everything but lions to pass through. This is then swept across the desert.
  2. 2.
    The Atom-Splitting Method. We irradiate the desert with slow neutrons. The lion becomes radioactive and a process of disintegration sets in. When the decay has proceeded sufficiently far, he will become incapable of showing fight.
  3. 3.
    The Magneto-Optical Method. We plant a large lenticular bed of catnip (Nepeta cataria), whose axis is aligned with the direction of the horizontal component of the earth’s magnetic field. We place a cage at one of its foci. The lion, attracted by the catnip, is bent towards the cage by the Faraday effect.
@@ 1.5 Earth as a Man-Controlled Spaceship
And now let’s look into a more distant future. Personally, I cannot understand why astronauts dream of getting into interstellar space. The main nuisance is that the astronaut in the rocket will be in the same position as a man placed against a beam of fast protons coming out of a powerful accelerator. I feel very sorry for the poor astronaut; I have even composed a ballad about his sad fate:
The Ballad of the Astronaut
The beta-inverters and gamma-converters Have stripped all the plating away, The ion-gun’s bust, crumbled to dust, And the nozzle is clogged with clay.
The mesons decayed, the neutrons have strayed, The light has all radiated. The protons have scattered (as if it mattered), And the leptons annihilated.
He flew to Orion, but a flux of gravitons Crossed his path in the dark. With resources spent and the hull badly bent, He barely missed the arc.
He fought gravity’s force, deviating from course, With super-acceleration, But the hands of the clock froze on the lock In time-dilation stagnation.
Back to the Earth! The land of his birth! But what is this red, cold ball? The sun has grown dim, the outlook is grim, There’s no one left to call.
With a curse and a cry at the frozen sky, He pulled the lever back, And "A" rang out, and "B" rang out, And then there was only… CRACK!
@@ 1.6 List of Typical Examination Questions for Physics Graduate Students
1.7 Introduction to S-Matrix Theory
(Considered mainly from the point of view of applications to the description of the life of physicists and primarily taking into account the statistical regularities characteristic of such systems) Original Source: The Journal of Irreproducible Results (1965).
It is well known that in recent years S-matrix theory has achieved significant success in describing scattering processes and the mutual transformation of elementary particles. This inspired us to attempt to apply it (perhaps not entirely rigorously) to the study of processes occurring with physicists throughout their lives. We will pay special attention to systems to which statistics can be applied, i.e., systems consisting of a large number of objects (in our case, physicists).
The system we are considering at time $t=$-$\infty$ represents an incident stream of physicists who can be considered almost free. According to two solutions of the equations of motion, this stream can be divided into two parts: retarded physicists and advanced physicists (the latter mainly from Princeton; they are distinguished by the fact that they never study the history of the question under consideration).
Throughout their lives, physicists enter into interaction with various systems. The strength of this interaction depends both on the skill and pushiness of each individual physicist, and on what these systems are—conservative or liberal. By the time $t=\infty$, the stream of physicists decays into various reaction products, the full number of which could in principle be obtained from known formulas for the S-matrix, if its form were currently known. The products can be distributed among so-called *reaction channels*, of which we will name here only a few:
  1. a)
    Scattered physicist;
  2. b)
    Professor;
  3. c)
    Mathematician;
  4. d)
    Reactor engineer;
  5. e)
    Bureaucrat.
From the most general properties of the S-matrix, and especially from its relativistic invariance, one can conclude that the total energy, including the rest mass, is an integral of the physicist’s motion through life. Since it is known that rest mass increases with age, we immediately conclude that the remaining energy falls over time.
To obtain more precise results, it is necessary to take into account the interaction of physicists with each other. For this purpose, let us consider a region of configuration space, the so-called "Institute," where the interaction is maximal. This region, hereafter called the CIPI (Configuration Interaction Physicist Interval) for brevity, is separated from the outside world by some potential barrier. Possible states of physicists in such a potential well can be defined by four quantum numbers, of which the first three have a generally known meaning. The fourth quantum number, corresponding to two possible states for a physicist—sleep and wakefulness—has no classical analogue, since, according to the quantum-mechanical principle of complementarity, neither of these states can be observed without an admixture of the other. We will denote the possible values of this quantum number by the symbols "+" and "–" respectively.
It is quite clear that the forces usually acting on physicists are so great that conducting any calculations using perturbation theory hardly seems expedient. Therefore, to obtain results, we must turn to simplified models. However, consideration of the latter would lead us far beyond the scope of this article. The results of these studies on models we will try to set out in subsequent works.
1.8 History of the Discovery of the Mössbauer Effect
(According to H. Lipkin) Original Source: Preface to H. Frauenfelder’s "The Mössbauer Effect".
Period Date Note
Prehistoric Up to 1958 Could have been discovered, but wasn’t.
Early Iridium Age 1958 Discovered, but not noticed.
Middle Iridium Age 1958–1959 Noticed, but not believed.
Late Iridium Age 1959 Believed, but… not interested.
Iron Age 1959–1960 OO–OO–OO–OO!!!
Dirac’s Puzzle
Dirac was distinguished by great ingenuity in solving various kinds of mathematical puzzles. In many cases, he offered his own, very unexpected solutions. A very popular puzzle is to express a given number using a limited number of identical digits, using any other mathematical signs. Dirac proposed a general solution to such a problem, finding a way to write any number using only three twos:
$N=-\log_{2}\log_{2}\sqrt{\sqrt{...\sqrt{2}}}$
The number of square root signs is equal to $N$.
Thermodynamics Summary
Among physicists, the following definition of thermodynamics is current:
"Thermodynamics is a stick with three ends."
@@ 1.9 Typology in Scientific Research
    THE DISCOVERER (The Idea Man). These scientists are always giving off sparks. Their brains are always ready to seize on random prey. They possess a high degree of scientific intuition, allowing them to formulate a hypothesis (or at least a working hypothesis) on the spot. They then either verify it experimentally themselves or, more likely, leave the drudgery of proof to others, deriving their pleasure from the purely speculative solution.
@@
    THE EXPLOITER. A researcher with a quick grasp; his eyes and ears are constantly open. This type is rarely found in his own laboratory; he prefers to spend his time in discussions with colleagues from other departments, especially if those colleagues are working on something interesting. He never lacks good ideas—usually borrowed from others—which he turns into interesting papers, generously sprinkled with references to "private communications."
@@
    THE CONNOISSEUR (The Critic). The mental abilities of this type far exceed his ability (or desire) to conduct his own experiments. He is capable of evaluating good work, often better than the author himself. However, his critical mind, combined with an innate inconsistency, prevents him from publishing anything, unless he has a decisive boss who forces the issue.
@@
    THE IMPROVER. He resembles the Connoisseur but possesses slightly higher productivity. His achievements are represented by very few, but excellent papers, based on experiments which he repeated so many times that all unexpected or unpredicted results were eventually discarded with the help of sophisticated statistical processing.
@@
    THE WELL-INFORMED (The Man Up-to-Date). He knows everything worth knowing. Unlike the Exploiter, he spends all his time in the library, where it is rare for anyone to beat him to the latest issue of a journal. He is a goldmine of references, but rarely produces original work, as he is too busy reading about the work of others.
@@
    THE CO-AUTHOR. This type has mastered the art of scientific diplomacy. He manages to include his name on the author list of most articles published by his department, often his only contribution being the decision of whether to use "and" or "&" in the title. Some hold the opinion that the "Co-author" is almost the same as the "Advisor," but everyone who plays bridge knows what a "dummy" is.
@@
    THE ADVISOR (The Consultant). Not to be confused with the Co-Author. The Advisor usually works for a rival institution or industry, giving advice to others for a fee, while doing no work of his own.
@@
    THE GADGETEER. In modern science, some research is impossible without a solid array of instruments. This type considers the meaning of life to be the acquisition of the largest possible number of ultra-modern instruments. His lab sparkles with glass and chrome; awe-inspiring names of installations are used to impress visitors rather than to obtain results.
@@
    THE PUBLISHER. This term denotes those who increase their list of publications at an exponential rate.
    • The Ruminant. (A name associated with the digestive process of certain mammals). This person usually settles in a developing country. He publishes his observations there, prefacing them with: "For the first time in the history of [Country] it was observed…", followed by a translation of work done years ago in another place.
    • The Multiplier. An individual who divides the results of his work into the largest possible number of small packets labeled "Article," "Letter to the Editor," "Preliminary Note," etc., successfully maximizing his bibliography.
@@
    THE CORRESPONDENT. To meet this type, one need only look at the "Letters" section of any journal. The Correspondent reports something "truly important," usually resembling a great discovery that must be published immediately to establish priority. Such messages end with: "A detailed description of these experiments will be published elsewhere." In 50% of cases, the promised publication never appears, as subsequent experiments destroy the interest in the idea.